BIGMONEYBEAR:

111-x^(3)-45^(4 )+205^(3)

BIGMONEYBEAR:

Find the x and y Intercepts

BIGMONEYBEAR:

x-intercept(s): (3√4514611,0) y-intercept(s): (0,4514611)

XioGonz:

`Step 1` 205 = 5*41 (205)^3 = (5*41)^3 = 5^3 * 41^3

BIGMONEYBEAR:

@tommymfng can u explain it tho

XioGonz:

`Which leaves us with` ((111 - (x^3)) - (45^4)) + (5^3•41^3)

XioGonz:

There is No Solution

BIGMONEYBEAR:

@xiogonz wrote:
`Which leaves us with` ((111 - (x^3)) - (45^4)) + (5^3•41^3)
good explanation

duhhitxmelivv:

111-x^3-45^4+205^3 45^4 = 4100625 111-x^3-4100625+205^3 205^3 = 8615125 111-x^3-4100625+8615125 Group Like terms: -x^3+111-4100625+8615125 Add subtract numbers: 111-4100625+ 8625125 =4514611 =-x^3+4514611 -x^3 + 4514611 x intercepts: (3^ 4514611,0) y intercept: (0, 4514611)